OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 2x + y = -4 5x + 3y = -6

OpenStudy (anonymous):

If we extend to first eq. with (-3) and sum to these equations we get (-x)=6 so x=-6 then if 2x+y=-4 y=8

OpenStudy (anonymous):

If we do this by elimination, we would multiply the top by -3, so the y-values would cancel out. That would give us -6x-3y=12 5x+3y=-6 Now we add/cancel them and get ---------- -x=6 [Divide both sides by -1] x=-6 Now that we know what the x-value is we will sub this back into any of the two equations and solve for y. I chose the 1st just because 2(-6)+y=-4 -12+y=-4 [Add 12 on both sides and get] y=8 That means our solution is (-6,8)